# The Great Ban: Efficacy and Unintended Consequences of a Massive Deplatforming Operation on Reddit

### v1

@@ Abstract
In the current landscape of online abuses and harms, effective content moderation is necessary to cultivate safe and inclusive online spaces. Yet, the effectiveness of many moderation interventions is still unclear. Here, we assess the effectiveness of The Great Ban, a massive deplatforming operation that affected nearly 2,000 communities on Reddit. By analyzing 16M comments posted by 17K users during 14 months, we provide nuanced results on the effects – both desired and otherwise – of the ban. Among our main findings is that 15.6% of the affected users left Reddit and that those who remained reduced their toxicity by 6.6% on average. The ban also caused 5% users to increase their toxicity by more than 70% of their pre-ban level. However, these resentful users likely had limited impact on Reddit due to low activity and little support by peers. Overall, our multifaceted results provide new insights into the efficacy of deplatforming. Our findings can inform the development of future moderation interventions and the policing of online platforms.
Introduction
@@
Within online platforms, content moderation is needed to maintain safe and inclusive social environments by mitigating the spread of problematic content and harmful behavior. It fosters user trust and safety, thereby upholding ethical standards and contributing to the overall flourishing of healthy online communities (Gillespie 2018). In content moderation, platform policies are enforced through moderation interventions (Trujillo, Fagni, and Cresci 2023). There exist many possible interventions to be applied by platform administrators, ranging from the delivery of short warning messages (Katsaros, Yang, and Fratamico 2022) and the use of informative labels (Zannettou 2021; Papakyriakopoulos and Goodman 2022), up to the removal of large amounts of content or users (Trujillo and Cresci 2022; Jhaver et al. 2021). However, in spite of the growing reliance on content moderation, there remains a limited understanding of the effects of many moderation interventions, which impair the efficacy of current regulatory efforts. Indeed, recent research has shown that some interventions yielded mixed (Horta Ribeiro et al. 2021; Trujillo and Cresci 2023) or no effects at all (Dias, Pennycook, and Rand 2020), and in some cases, some even resulted in unintended and undesired consequences (Bail et al. 2018; Pennycook et al. 2020). For these reasons, it is paramount to conduct thorough assessments of the outcomes of recent moderation interventions as a fundamental preliminary step to the planning and development of future effective solutions.
@@
Out of all the possible interventions, the removal (i.e., banning) of users, communities, and content – a moderation practice dubbed deplatforming – is by far the most frequently adopted (Horta Ribeiro et al. 2024). A famous example is the banishment of Donald Trump from Facebook and X (formerly Twitter) in 2021 (Seeliger and Baum 2023). Other notable examples are the deplatforming of some toxic influencers on X (Jhaver et al. 2021) and the permanent shut of racist, sexist, and generally hateful communities on Reddit (Chandrasekharan et al. 2017; Habib et al. 2022). In 2020 Reddit hosted a massive deplatforming operation that involved the ban of around 2,000 subreddits on the grounds of their promotion of identity- and vulnerability-based hate. [^1] https://www.reddit.com/r/announcements/comments/hi3oht/update˙to˙our˙content˙policy/ (accessed: 12/15/2023) Among the banned communities were some very popular subreddits such as r/The_Donald and r/ChapoTrapHouse. This event is commonly referred to as The Great Ban and is currently one of the biggest bans in the history of social media. Despite its impact on multiple communities and on a large number of users within and beyond Reddit, its effects remain underexplored. Existing studies on The Great Ban have primarily focused on assessing writing style changes (Trujillo et al. 2021), which do not answer the fundamental question as to whether the ban was effective at curbing toxic behavior. The few studies that sought to analyze the degree of toxicity after the ban did so only for a few of the affected subreddits (Trujillo and Cresci 2022). Moreover, the majority of such studies investigated community-level effects, overlooking user-level reactions that are however clearly relevant with respect to the moderation goal of thwarting particularly problematic behaviors (Trujillo and Cresci 2023).
@@
  • However, their impact was likely minimal due to low activity levels and little support from other Reddit users.
Overall, our study provides a comprehensive account of the effects of The Great Ban. It surfaces and describes an undesired side effect of the intervention, drawing attention on the delicate balance entailed by the moderation of heterogeneous communities. As such, our results can inform future moderation strategies and the development of effective interventions.
Related Work
@@
Deplatforming
Despite the relevance and the extent of The Great Ban, few works delved into a systematic evaluation of its effects. Among them is the study by Milo Trujillo et al. (2021) that analyzed activity and linguistic changes in the 15 most popular subreddits affected by the ban. They found that top users suffered the largest decreases in activity and that community response was heterogeneous between subreddits, and even between users of a subreddit (Trujillo et al. 2021). Here we complement this work by assessing effects in terms of toxicity, rather than activity and language. Other works evaluated deplatforming effects in a subset of the subreddits affected by The Great Ban, or in other subreddits altogether. Chandrasekharan et al. (2017) and Saleem and Ruths (2018) examined the repercussions of the bans on r/fatpeoplehate and r/coontown, revealing that a substantial number of users departed from Reddit following the interventions. Among those who stayed, a notable reduction in hate speech was observed. However, they also found a considerable portion of users who migrated to other subreddits, doubling their posting activity (Chandrasekharan et al. 2022). Instead, Horta Ribeiro et al. (2021) assessed the impact of deplatforming across multiple platforms, concentrating on the migration of users from banned subreddits to newly established platforms. Their findings indicated a significant decline in user activity on the new platforms. However, they also found a subset of users who increased their toxicity and radicalization. Deplatforming was also studied on platforms other than Reddit. For example, Mekacher, Falkenberg, and Baronchelli (2023) studied ban-induced migrations from Twitter to Gettr, finding that politically polarized users are less toxic on fringe platforms as they are exposed to less out-group interactions. Finally, Jhaver et al. (2021) examined Twitter’s ban on multiple toxic influencers, revealing a general reduction in conversations about these figures, accompanied by decreased activity and toxicity among their supporters. Nonetheless, they also found a fraction of users who greatly increased activity and toxicity.
Overall, this body of work shows that moderation interventions frequently cause a combination of desired and undesired effects, and that effects vary between different interventions and types of users. To this end, our work leverages and extends previous knowledge by evaluating the effects of The Great Ban – a massive, yet essentially unexplored, moderation event – across the main dimensions of user behavior: toxicity, activity, and engagement. Our results provide a complete picture of the effectiveness of the ban on the 15 most popular affected subreddits.
@@
Table 1: Dataset composition. Rows are ordered by number of active users after the ban. Data in IN-BEFORE is related to user activities within the banned subreddits before the ban took place. Data in OUT-BEFORE and OUT-AFTER are related to user activities outside of the banned subreddits, respectively before and after the ban. OUT-BEFORE contains 288 less users than IN-BEFORE, who had no activity outside of the banned subreddits. OUT-AFTER contains 2,577 less users than OUT-BEFORE, who abandoned the platform after The Great Ban.
Soft moderation
So-called soft interventions emerged as an alternative to deplatforming, addressing the concerns about censorship and the loss of free speech that often accompany content and user removals (Zannettou 2021). Amaury Trujillo and Cresci (2022; 2023) studied quarantines and restrictions: soft interventions that often precede community bans on Reddit. They studied moderation outcomes on r/the_donald, revealing a general reduction in activity and toxicity, at the cost of an increased political polarization and decreased factuality of shared news (Trujillo and Cresci 2022, 2023). Reddit’s quarantine of r/the_donald was also studied by Chandrasekharan et al. (2022) and Shen and Rosé (2022), who concluded that the intervention did not produce meaningful changes, nor in terms of misogyny and racist comments, neither regarding engagement and internal dynamics. Other types of soft interventions are the attachment of warning labels to disputed posts, whose effects were assessed in terms of perceived credibility and obtained engagement. In detail, Pennycook et al. (2020) found that the presence of some posts with warning labels increases the perceived credibility of all posts without labels, including false ones not yet debunked. Zannettou (2021) found instead that tweets with warning labels were often replied to, to further debunk the claims. However, this resulted in the flagged tweets circulating more and obtaining more engagement than the undisputed ones. Finally, Katsaros, Yang, and Fratamico (2022) carried out an A/B test on Twitter to evaluate the effectiveness of using warning messages to prompt users that are about to post toxic tweets. Their results show that the intervention was overall effective at reducing the posting of toxic tweets. Nonetheless, a small minority of users edited their tweet to make it more toxic after being exposed to the warning message.
@@
Figure 1: Timeline depicting the periods of data collection and analysis. Collected data spans two time periods of 7 months each, centered around The Great Ban. It is noteworthy that the IN-BEFORE dataset has no content since May 2020, indicating that the activity in the banned subreddits halted before the official intervention date.
Dataset
@@
Our dataset for this study comprises 16M Reddit comments shared by 16,828 distinct users who participated in at least one of the 15 popular subreddits shut during The Great Ban (Trujillo et al. 2021), as reported in Table 1. The composition of our dataset is illustrated in Figure 1 and the procedure adopted to build it is described in the following.
@@
Dataset construction. We initially collected all comments posted between December 2019 and June 2020 in each of the 15 popular subreddits, resulting in 8M comments shared by 194K distinct users. The data covers 30 weeks (i.e., 7 months) prior to The Great Ban, and allows for establishing a suitable baseline for the activity of the affected users before the moderation intervention (Trujillo and Cresci 2023). Notably, we could not collect any data between May and June 2020, due to several subreddits having halted their activity, or being banned, before Reddit’s public announcement of The Great Ban on June 29, 2020.
@@
As commonly done in literature, we obtained a representative set of users for the considered subreddits by constraining our analysis to core users – namely, users who participated regularly in at least one subreddit (Trujillo and Cresci 2022; Bouleimen et al. 2023). We defined core users as those who posted at least one comment each month between December 2019 and March 2020.222April 2020 was excluded from this analysis due to the limited number of collected comments. It is likely that the activity within the subreddits halted in the initial days of that month. Additionally, we filtered out bots (i.e., clearly automated accounts) by discarding all accounts that posted at least two different comments at the exact same time. After these filtering steps, we ended up with 2.2M comments by 16,828 core users. Hereafter, we refer to this portion of our dataset as IN-BEFORE since it involves activity within the banned subreddits before the ban.
@@
Providing a fair evaluation of the effects of The Great Ban involves matching comparable datasets before and after the intervention. However, no activity exists within the banned subreddits after the intervention, since the subreddits were permanently shut. Therefore, evaluating the effects of the intervention must involve the analysis of user activities outside of the banned subreddits. For this reason, we collected all comments made by the 16,828 core users outside of the 15 banned subreddits across a wide time frame spanning 7 months before and after The Great Ban, as shown in Figure 1. We obtained around 13.8M comments by 16,540 distinct users. We labeled data related to user activities before the ban as OUT-BEFORE and that after the ban as OUT-AFTER, as reported in Table 1. Estimates of the effects of the ban are obtained by comparing the OUT-BEFORE and OUT-AFTER datasets. Lastly, we enriched our dataset by computing a toxicity score for each collected comment.
@@
Figure 2: Comparison between the toxicity (left panel) and severe toxicity (right panel) indicators provided by Perspective API (x axis) and Detoxify (y axis). Overall, we report a strong Pearson correlation between the two methods, and particularly so for the toxicity indicator ($r=0.872$).
@@
Annotating toxicity: Detoxify vs Perspective API. Google’s Perspective API and Detoxify are among the state-of-the-art and most widely used methods for computing toxicity scores for texts. Perspective API [^3] https://perspectiveapi.com/ (accessed: 01/01/2024) was developed by the Jigsaw team at Google and currently represents the de facto standard for toxicity detection, both in production content moderation environments and in academia (Rieder and Skop 2021; Nogara et al. 2023). The service is offered as a Web API that, given a piece of text, outputs several scores of offensiveness, including two indicators of toxicity and severe toxicity defined in the [0, 1] range. Detoxify is an open source deep learning toxicity classifier that also outputs the toxicity and severe toxicity indicators (Hanu and Unitary team 2020). Due to its convenience and competitive performance, it has recently seen frequent use (Chang, May, and Lerman 2023; Noor and Ahmed 2023). The advantage of Detoxify over Perspective API lies in the possibility to install and run it locally, without incurring in the limitations of a Web API (i.e., rate limits or quotas). Given the large number of comments to annotate in our dataset, we computed toxicity scores with Detoxify. Nonetheless, we first validated our choice by comparing the outputs of Perspective API and Detoxify on a stratified random sample of 10K comments extracted from the IN-BEFORE portion of our dataset. Figure 2 presents the results of this comparison, for both the toxicity and severe toxicity indicators. As shown, we found a strong positive Pearson correlation between the two methods, which supports the use of Detoxify. Then, with respect to the two provided indicators, for our subsequent analyses we relied on the toxicity indicator because it is the one on which the two methods agree the most, with $r=0.872$ vs $r=0.796$ for severe toxicity.
Analyses and Results

### v5

@@ Abstract.
In the current landscape of online abuses and harms, effective content moderation is necessary to cultivate safe and inclusive online spaces. Yet, the effectiveness of many moderation interventions is still unclear. Here, we assess the effectiveness of The Great Ban, a massive deplatforming operation that affected nearly 2,000 communities on Reddit. By analyzing 16M comments posted by 17K users during 14 months, we provide nuanced results on the effects —both desired and otherwise— of the ban. Among our main findings is that 15.6% of the affected users left Reddit and that those who remained reduced their toxicity by 6.6% on average. The ban also caused 5% users to increase their toxicity by more than 70% of their pre-ban level. Overall, our multifaceted results provide new insights into the efficacy of deplatforming. As such, our findings can inform the development of future moderation interventions and the policing of online platforms.
Content moderation, toxicity, deplatforming, Reddit
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Diffusion of Harmful Content on Online Web Workshop; May 21-24, 2024; Stuttgart, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computingccs: Information systems Social networks
1. Introduction
@@
Within online platforms, content moderation is needed to maintain safe and inclusive social environments by mitigating the spread of problematic content and harmful behavior. It fosters user trust and safety, thereby upholding ethical standards and contributing to the overall flourishing of healthy online communities (Gillespie, 2018). Hence, platform policies are enforced through moderation interventions (Trujillo et al., 2023). There exist many possible interventions to be applied by a platform, ranging from short warning messages (Katsaros et al., 2022) and the use of informative labels (Zannettou, 2021; Papakyriakopoulos and Goodman, 2022) up to the removal of large amounts of content or users (Trujillo and Cresci, 2022; Jhaver et al., 2021). However, in spite of the growing reliance on content moderation, there is still a limited understanding of the general effectiveness of moderation interventions, which impair the efficacy of current regulatory efforts. Indeed, recent research has shown that some interventions yielded mixed (Horta Ribeiro et al., 2021; Trujillo and Cresci, 2023) or no effects at all (Dias et al., 2020), and in some cases, some even resulted in unintended and undesired consequences (Bail et al., 2018; Pennycook et al., 2020). For these reasons, it is paramount to conduct thorough assessments of the outcomes of recent moderation interventions as a fundamental preliminary step to the planning and development of future effective solutions.
@@
Out of all the possible interventions, the removal (i.e., banning) of users, communities, and content —a moderation practice called deplatforming— is by far the most frequently adopted (Horta Ribeiro et al., 2024). A famous example is the banishment of Donald Trump from Facebook and X (formerly Twitter) in 2021 (Seeliger and Baum, 2023). Other notable examples are the deplatforming of some toxic influencers on X (Jhaver et al., 2021), the removal of accounts involved in coordinated inauthentic behaviors on X (Cima et al., 2024), and the permanent shut of racist, sexist, and generally hateful communities on Reddit (Chandrasekharan et al., 2017; Habib et al., 2022). In 2020, Reddit carried out a massive deplatforming operation that involved the ban of around 2,000 subreddits, on the grounds of their promotion of hate of groups based on identity and vulnerability. [^1] https://www.reddit.com/r/announcements/comments/hi3oht/update_to_our_content_policy/ (accessed: 03/15/2024) Among the banned communities were the very popular subreddits r/The_Donald and r/ChapoTrapHouse, with hundreds of thousands of subscribers. This event is commonly referred to as The Great Ban and is currently one of the biggest bans in the history of social media. Despite its impact on multiple communities and on a large number of users within and beyond Reddit, its effects remain under-explored. Existing studies on The Great Ban have primarily focused on assessing writing style changes (Trujillo et al., 2021), which do not answer the fundamental question as to whether the ban was effective at curbing toxic behavior. The few studies that sought to analyze the degree of toxicity after the ban did so only for a few of the affected subreddits (Trujillo and Cresci, 2022). Moreover, the majority of such studies investigated community-level effects, overlooking user-level reactions that are however clearly relevant with respect to the moderation goal of thwarting particularly problematic users and behaviors (Trujillo and Cresci, 2023).
@@
Overall, our study provides a comprehensive account of the effects of The Great Ban. It surfaces and describes an undesired side effect of the intervention, drawing attention to the delicate balance entailed by the moderation of heterogeneous communities. As such, our results can inform future moderation strategies and the development of effective interventions.
2. Related Work
@@
2.1. Deplatforming
Despite the relevance and the extent of The Great Ban, few works delved into a systematic evaluation of its effects. Among them is the study by Milo Trujillo et al. ((2021)) that analyzed activity and linguistic changes in the 15 most popular subreddits affected by the ban. They found that top users suffered the largest decreases in activity and that community response was heterogeneous between subreddits, and even between users of a subreddit (Trujillo et al., 2021). Here we complement this work by assessing effects in terms of toxicity, rather than activity and language. Other works evaluated deplatforming effects in a subset of the subreddits affected by The Great Ban, or in other subreddits altogether. Chandrasekharan et al. (2017) and Saleem and Ruths (2018) examined the repercussions of the bans on r/fatpeoplehate and r/coontown, revealing that a substantial number of users departed from Reddit following the interventions. Among those who stayed, a notable reduction in hate speech was observed. However, they also found a considerable portion of users who migrated to other subreddits, doubling their posting activity (Chandrasekharan et al., 2022). Instead, Horta Ribeiro et al. (2021) assessed the impact of deplatforming across multiple platforms, concentrating on the migration of users from banned subreddits to newly established platforms. Their findings indicated a significant decline in user activity on the new platforms. However, they also found a subset of users who increased their toxicity and radicalization. Deplatforming was also studied on platforms other than Reddit. For example, Mekacher et al. (2023) studied ban-induced migrations from Twitter to Gettr, finding that politically polarized users are less toxic on fringe platforms as they are exposed to less out-group interactions. Cima et al. (2024) analyzed massive bans done by Twitter to counteract coordinated inauthentic behaviors. Finally, Jhaver et al. (2021) examined Twitter’s ban on multiple toxic influencers, revealing a general reduction in conversations about these figures, accompanied by decreased activity and toxicity among their supporters. Nonetheless, they also found a fraction of users who greatly increased activity and toxicity.
Overall, this body of work shows that moderation interventions frequently cause a combination of desired and undesired effects, and that effects vary between different interventions and types of users. To this end, our work leverages and extends previous knowledge by evaluating the effects of The Great Ban —a massive, yet essentially unexplored, moderation event— across the dimension of comment toxicity. Our results provide a picture of the effectiveness of the ban on the 15 most popular affected subreddits.
@@ 2.1. Deplatforming
Table 1. Dataset composition. Rows are ordered by number of active users after the ban. Data in IN-BEFORE is related to user activities within the banned subreddits before the ban took place. Data in OUT-BEFORE and OUT-AFTER are related to user activities outside of the banned subreddits, respectively before and after the ban. OUT-BEFORE contains 288 less users than IN-BEFORE, who had no activity outside of the banned subreddits. OUT-AFTER contains 2,577 less users than OUT-BEFORE, who abandoned the platform after The Great Ban.
2.2. Soft moderation
So-called soft interventions emerged as an alternative to deplatforming, addressing the concerns about censorship and the loss of free speech that often accompany content and user removals (Zannettou, 2021). Trujillo and Cresci ((2022; 2023)) studied quarantines and restrictions: soft interventions that often precede community bans on Reddit. They studied moderation outcomes on r/the_donald, revealing a general reduction in activity and toxicity, at the cost of an increased political polarization and decreased factuality of shared news (Trujillo and Cresci, 2022, 2023). Reddit’s quarantine of r/the_donald was also studied by Chandrasekharan et al. (2022) and Shen and Rosé (2022), who concluded that the intervention did not produce meaningful changes, nor in terms of misogyny and racist comments, neither regarding engagement and internal dynamics. Another type of soft intervention is the attachment of warning labels to disputed posts, whose effects were assessed in terms of perceived credibility and obtained engagement. In detail, Pennycook et al. (2020) found that the presence of some posts with warning labels increases the perceived credibility of all posts without labels, including false ones not yet debunked. Zannettou (2021) found instead that tweets with warning labels were often replied to, to further debunk the claims. However, this resulted in the flagged tweets circulating more and obtaining more engagement than the undisputed ones. Finally, Katsaros et al. (2022) carried out an A/B test on Twitter to evaluate the effectiveness of using warning messages to prompt users who are about to post toxic tweets. Their results show that the intervention was overall effective at reducing the posting of toxic tweets. Nonetheless, a small minority of users edited their tweets to make them more toxic after being exposed to the warning message.
@@ 2.2. Soft moderation
Figure 1. Timeline depicting the periods of data collection and analysis. Collected data spans two time periods of 7 months each, centered around The Great Ban. It is noteworthy that the IN-BEFORE dataset has no content since May 2020, indicating that the activity in the banned subreddits halted before the official intervention date.
3. Dataset
@@
Our dataset for this study comprises 16M Reddit comments shared by 16,828 distinct users who participated in at least one of the 15 most popular public subreddits in terms of daily active users shut during The Great Ban (Trujillo et al., 2021), as reported in Table 1. Even if Reddit administrators originally published an obfuscated list of the most popular banned subreddits, [^2] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 03/15/2024) previous work deciphered it for those with over 2,000 daily active users (Trujillo et al., 2021), resulting in the 15 subreddits used both therein and herein. The composition of our dataset is illustrated in Figure 1 and the procedure adopted to build it is described in the following.
@@
Dataset construction. We initially collected all comments posted between December 2019 and June 2020 in each of the 15 popular subreddits, resulting in 8M comments shared by 194K distinct users. To collect the comments we used the Pushshift data dumps (Baumgartner et al., 2020) available through Reddit torrents. [^3] https://academictorrents.com/details/9c263fc85366c1ef8f5bb9da0203f4c8c8db75f4 (accessed: 03/15/2024) The data covers 30 weeks (i.e., 7 months) prior to The Great Ban, which allows for establishing a suitable baseline for the activity of the affected users before the moderation intervention (Trujillo and Cresci, 2023). Notably, we could not collect any data between May and June 2020, due to several subreddits having halted their activity, or being banned, before Reddit’s public announcement of The Great Ban on June 29, 2020.
@@
As commonly done in literature, we obtained a representative set of users for the considered subreddits by constraining our analysis to core users —namely, users who participated regularly in at least one subreddit (Trujillo and Cresci, 2022; Bouleimen et al., 2023). We defined core users as those who posted at least one comment each month between December 2019 and March 2020.444April 2020 was excluded from this analysis due to the limited number of collected comments. It is likely that the activity within the subreddits halted in the initial days of that month. Additionally, we filtered out bots (i.e., clearly automated accounts) by discarding all accounts that posted at least two different comments at the exact same time. For this, we used a time delta of at least 1 second between comments, which guarantees we do not inadvertently filter out authentic users (Hurtado et al., 2019). Additionally, we manually validated a random sample of 1,000 comments to verify that manifest bots were effectively excluded from the dataset. After these filtering steps, we ended up with 2.2M comments by 16,828 core users. Hereafter, we refer to this portion of our dataset as IN-BEFORE since it involves activity within the banned subreddits before the ban.
@@
Providing a fair evaluation of the effects of The Great Ban involves matching comparable datasets before and after the intervention. However, no activity exists within the banned subreddits after the intervention, since the subreddits were permanently shut. Therefore, evaluating the effects of the intervention must involve the analysis of user activities outside of the banned subreddits. For this reason, we collected all comments made by the 16,828 core users outside of the 15 banned subreddits across a wide time frame spanning 7 months before and after The Great Ban, as shown in Figure 1. We obtained around 13.8M comments from 16,540 distinct users. We labeled data related to user activities before the ban as OUT-BEFORE and that after the ban as OUT-AFTER, as reported in Table 1. Estimates of the effects of the ban are obtained by comparing the OUT-BEFORE and OUT-AFTER datasets. Lastly, we enriched our dataset by computing a toxicity score for each collected comment.
@@
Figure 2. Comparison between the toxicity (left panel) and severe toxicity (right panel) indicators provided by Perspective API (x axis) and Detoxify (y axis). Overall, we report a strong Pearson correlation between the two methods, and particularly so for the toxicity indicator ($r=0.872$).
@@
Annotating toxicity: Detoxify vs Perspective API. Google’s Perspective API and Detoxify are among the state-of-the-art and most widely used methods for computing toxicity scores for texts. Perspective API [^5] https://perspectiveapi.com/ (accessed: 03/15/2024) was developed by the Jigsaw team at Google and currently represents the de facto standard for toxicity detection, both in production content moderation environments and in academia (Rieder and Skop, 2021; Nogara et al., 2023). The service is offered as a Web API that, given a piece of text, outputs several scores of offensiveness, including two indicators of toxicity and severe toxicity defined in the [0, 1] range. Detoxify is an open source deep learning toxicity classifier that also outputs the toxicity and severe toxicity indicators (Hanu and Unitary team, 2020). Due to its convenience and competitive performance, it has recently seen frequent use (Ejaz et al., 2024; Köpf et al., 2024). The advantage of Detoxify over Perspective API lies in the possibility of installing and running it locally, without incurring the limitations of a Web API (i.e., rate limits or quotas). Given the large number of comments to annotate in our dataset, we computed toxicity scores with Detoxify. Nonetheless, we first validated our choice by comparing the outputs of Perspective API and Detoxify on a stratified random sample of 10K comments extracted from the IN-BEFORE portion of our dataset. Figure 2 presents the results of this comparison, for both the toxicity and severe toxicity indicators. As shown, we found a strong positive Pearson correlation between the two methods, which supports the use of Detoxify. Then, with respect to the two provided indicators, for our subsequent analyses we relied on the toxicity indicator because it is the one on which the two methods agree the most, with $r=0.872$ vs $r=0.796$ for severe toxicity.
4. Analyses and Results